A tugboat goes 180 miles up stream in 15 hours. The return trip down stream takes 10 hours. Find the speed of the tugboat with out a currant and the speed of the currant.
The speed of the tugboat is __ mph and the speed of the currant is __ mph
simplify your answers

Answer by richwmiller(17219)

R*T=D
R*15=180
R=12
r*10=180
r=18
18-12=6
6/2=3
stream = 3 mph
tugboat goes 15 mph without current.